OpenStudy (anonymous):

The sum of two numbers is 4 less than twice the larger. If the larger is decreased by three times the smaller, the result is -20. Find the numbers

OpenStudy (anonymous):

16,12 ?

OpenStudy (anonymous):

Here i are the equations: x is the larger and y is the smaller. x + y = 2x-4 x -3y = -20. solve the above syatem of equations to obtain the values of x and y

OpenStudy (anonymous):

Rearange the equations as follows: -x + y = -4 ( after subtracting 2x from both sides) x -3y = -20 add both equations to eliminate x i.e. -2y = -24 y =12 Plug in the value of y in any one of th above equations to yield -x +12 = -4 =>-x = -16 => x = 16 Therefore larger number = 16 Smaller number =12

OpenStudy (anonymous):

You can verify the results by walking through the quation using the above numbers. does this solve your problem?